Exercise 5.10
A cylinder with a moveable piston contains 0.91 mol of gas and has a volume of 315 mL .

What is its volume after an additional 0.16 mol of gas is added to the cylinder? (Assume constant temperature and pressure.)
Express your answer to three significant figures.
1
Expert's answer
2017-03-24T08:52:15-0400
Answer:
pV=nRT
p·0.000315=0.91·RT
In order the balance to be maintained, equal correlation must be in place. In this case correlation between the figures is: 0.91/0.000315 = 2888.889
n1 = 0.91+0.16 = 1.07 mol
p ·X = 1.07 ·RT
х = 1.07/2888.889 = 0.000370 m3 = 370 ml